There were 52 questions on a test. there were three times as many true and false questions as multiple choice questions. solve a system to find the number of each type of questions.

Found 2 solutions by checkley79, Ashu_321:
Answer by checkley79(3341)

M+T=52
T=3M
M+3M=52
4M=52
M=52/4
M=13 ANS. FOR THGE NUMBER OF MULTIPLE CHOICE QUESTIONS.
T=3*13=39 ANS. FOR THE TRUE FALSE QUESTIONS.
PROOF:
13+39=52
52=52

Answer by Ashu_321(1)